OpenStudy (anonymous):

what is 6\49 *17\48 *8\3 *6 multiply

OpenStudy (anonymous):

(6/49)(17/48) = 102/2352 (102/2352)(8/3) = 816/7056 (816/7056)(6/1) = 4896/7056

OpenStudy (anonymous):

in simpliest form

OpenStudy (anonymous):

first try taking a common even number out... say 16. so that simplifies to 306/441 then take i tried 9 so 34/49.

OpenStudy (anonymous):

\[(6/49)*(17/48)*(8/3)*6 = \frac{34}{49} \]